What is the probability of picking an ace from a pack of playing cards?
The probability of drawing an Ace from a standard deck of 52 cards is 4 in 52, or 1 in 13, or about 0.07692.
What is the probability of getting court card?
The probability of drawing a court card (a jack, queen, or king) from a standard deck of 52 cards is 12 in 52, or 3 in 13, or about 0.2308.
If you count the ace as a court card, then the probability is 16 in 52, or 4 in 13, or about 0.3077.
When a single card is drawn from an ordinary 52-card deck find the probability of getting a diamond?
There are 13 diamonds in a 52-card deck
so the probability to get a diamond is 13/52=1/4=0.25
What is the probability of drawing a black card from a 52 card deck and flipping heads on a coin?
Half the cards in a standard pack are black.
Therefore the probability of drawing a black card is 1/2.
Half the sides of a coin are "heads" so again the probability is 1/2.
Therefore the probability you will both draw a black card and flip heads = 1/2 * 1/2 = 1/4.
What is the probability of being dealt 3 cards and all of them being diamonds?
There are 13 diamonds. Three cards are dealt. The probability of all of them being diamond is (13/52)(12/51)(11/50) = 1716/132600 = 11/850
No, a probability is a number between 0 (included) and 1 (included)
What is the probability it will rain our of four days?
The probability is 1.
It is always raining - any time of the day or night, any day of the year - somewhere on Earth.
What is required in a binomial distribution?
A number of trials, each of which has only two outcomes: these are usually termed "success" and "failure". The trials must be independent and the probability of success must remain constant.
What is the probability of spinning an odd number 8 1 7 2 6 5 4 3?
8 numbers in total, 4 of those numbers are odd (1, 7, 5 and 3)
so 4/8
= 1/2
aka 50%, 0.5
How many permutations are in the word math?
Since the word MATH does not have any duplicated letters, the number of permutations of those letters is simply the number of permutations of 4 things taken 4 at a time, or 4 factorial, or 24.
What is the variance of the uniform distribution function?
the variance of the uniform distribution function is 1/12(square of(b-a)) and the mean is 1/2(a+b).
How many marbles fit into an empty paper bag?
It depends on the size of the marbles and on the size of the paper bag, neither of which was given in the question. Please restate the question.
How many distinct permutations are there of the word statistics?
We know there are 10! (ten factorial) permutations (that's about 3,628,800 permutations); however, we know that number includes repeated permutations, as there are 3 s', 3 t's and 2 i's. So we have to divide by the number of ways these can be written as individual permutations (if they were considered as unique elements), which are 3! (= 6), 3! and 2! (= 2) respectively.
So our final calculation would be 3628800 / (6 * 6 * 2) = 50400 unique permutations.
When to use cumulative binomial probability?
When the event of interest is a cumulative event.
For example, to find the probability of getting three Heads in 8 tosses of a fair coin you would use the regular binomial distribution. But to find the probability of up to 3 Heads you would use the cumulative distribution.
This is because
Prob("up to 3") = Prob(0 or 1 or 2 or 3) = Prob(0) + Prob(1) + Prob(2) + Prob(3) since these are mutually exclusive.
What does likely mean in math?
Likely means the same as probable and probable refers to the probability that something might happen,
Probability is measured using a scale of probability from zero to one.
Zero means that the event will definitely not happen whereas a probability of 1 means that the event is certain to happen, Similarly a probability of 0.5 means that the event is just as likely not to happen as to happen.
What is the probability that you could swim the Atlantic unaided? Zero
What is the probability that a bird will die? One
What is the probability that I can throw a 1 with a dice? 1 out of 6 = 0.1666666.......
What is the probability that I can toss a "heads" with a coin? 50% = 0.5 = one half
What is the probability of drawing a red two?
The probability of drawing a red two from a standard deck of 52 cards is 2 in 52, or about 0.03846.
What is the probability of getting an odd number when rolling a single 6-sided dice?
The probability of rolling an odd number of a standard die is 3 in 6, or 1 in 2, or 0.5.
What is the probability of being dealt a club out of a standard deck of cards?
There are 13 clubs in a standard deck of 52 cards. The probability, then, of drawing club is 13 in 52, or 1 in 4, or 0.25.
A card is drawn from a standard deck of playing cards. what is the probability that a spade and a heart is selected?
There are 2 of the threes that meet your condition.
If rolling three dice what is the probability of rolling three ones?
The probability of of rolling three ones on three dice is (1 in 6)3, or 1 in 216, or about 0.004630.
How do you calculate the third quartile for a mean of 110 and a standard deviation of 15?
If this is the only information you have, the answer would be somewhere around 125. Usually, you would find the third quartile by first finding the median. Then find the median of all of the numbers between the median and the largest number, which is the third quartile.
From a deck of 52 playing cards what is the probability that the card is a queen or a red card?
In a standard deck of 52 cards, there are 13 red cards and four queens. Since one of the red cards is also a queen, there are 16 cards that are either red or a queen. The probability, then, of drawing a red card or a queen is 16 in 52, or 8 in 26, or 4 in 13, or about 0.3077.
What is the probability of selecting a face card in a deck of 52 cards?
There are 12 face cards in a standard deck of 52 cards; the jacks, queens, and kings of spades, diamonds, clubs, and hearts. The probability, then, of drawing a face card is 12 in 52, or 3 in 13, or about 0.2308.